# Robust Regularized Policy Iteration under Transition Uncertainty

### 10 Mar 2026 v1

@@ Abstract
Offline reinforcement learning (RL) enables data-efficient and safe policy learning without online exploration, but its performance often degrades under distribution shift. The learned policy may visit out-of-distribution state-action pairs where value estimates and learned dynamics are unreliable. To address policy-induced extrapolation and transition uncertainty in a unified framework, we formulate offline RL as robust policy optimization, treating the transition kernel as a decision variable within an uncertainty set and optimizing the policy against the worst-case dynamics. We propose Robust Regularized Policy Iteration (RRPI), which replaces the intractable max-min bilevel objective with a tractable KL-regularized surrogate and derives an efficient policy iteration procedure based on a robust regularized Bellman operator. We provide theoretical guarantees by showing that the proposed operator is a $\gamma$-contraction and that iteratively updating the surrogate yields monotonic improvement of the original robust objective with convergence. Experiments on D4RL benchmarks demonstrate that RRPI achieves strong average performance, outperforming recent baselines including percentile-based methods such as PMDB on the majority of environments while remaining competitive on the rest. Moreover, RRPI exhibits robust behavior. The learned $Q$-values decrease in regions with higher epistemic uncertainty, suggesting that the resulting policy avoids unreliable out-of-distribution actions under transition uncertainty.
@@ 1 Introduction
Reinforcement learning (RL) has become a key ingredient of modern decision-making systems, driving breakthroughs in complex domains such as strategic game playing, embodied intelligence, and large language model reasoning by optimizing long-horizon policies from interaction data [26, 15, 34, 27]. Despite these successes, standard online RL typically relies on trial-and-error exploration, a process that can be prohibitively costly in high-stakes real-world applications [5, 9]. This has driven growing interest in offline RL [11], which aims to learn high-performing policies solely from pre-collected datasets without additional environment interaction.
@@
In this work, we pursue a robust policy optimization perspective [20]. Instead of treating the learned dynamics as a point estimate, we regard the transition model as a decision variable that lies in an uncertainty set of plausible kernels, and we seek a policy that maximizes performance under the worst-case dynamics in this set [1]. This formulation incorporates dynamics uncertainty directly into the objective and yields a principled alternative to heuristic uncertainty penalties. Nevertheless, the induced max-min objective leads to a challenging bilevel optimization problem, whose exact solution is often computationally prohibitive in practical offline RL pipelines. To address this, we propose Robust Regularized Policy Iteration (RRPI), which introduces a KL-regularized surrogate objective and an associated robust regularized Bellman operator, enabling an efficient iterative algorithm with theoretical guarantees.
@@ 2.1 Related work
Offline RL learns a policy solely from a fixed dataset, and its core challenge arises from distributional shift: the learned policy may query state-action pairs that are poorly covered by the data, making value estimates unreliable [11, 21] and leading to extrapolation error [3, 18]. Uncertainty in offline RL is often characterized as epistemic uncertainty induced by limited data support, which differs from aleatoric uncertainty originating from stochastic environment dynamics and observation noise [6, 4]. To mitigate the epistemic uncertainty, some efforts adopt conservative value learning that penalizes OOD actions [10, 32, 7, 31], regularized policy updates that constrain deviations from the behavior policy [29, 39, 37], and explicit uncertainty estimation via ensembles or Bayesian methods [23, 6, 22]. In contrast to prior methods, we cast offline RL as robust optimization over an uncertainty set, which incorporates dynamics uncertainty directly without explicit uncertainty estimation. In this formulation, the dynamics model is treated as a variable rather than a fixed quantity. This method avoids heuristic penalties on OOD action values and does not require constraining the learned policy to remain close to the behavior policy. Building on this, we develop an iterative policy optimization algorithm with theoretical convergence guarantees to an optimal solution.
@@ 2.2 Background
The modeling power and conservatism of robust MDPs depend critically on how $P$ is constructed. A common assumption is rectangularity, where $P$ factorizes across states or state-action pairs, enabling the inner minimization to decompose locally and making robust dynamic programming tractable. While convenient, rectangular sets can be overly conservative in high-dimensional problems since they allow independent worst-case perturbations at each $(s,a)$. In practice, $P$ is often built from function approximation, for example by using model ensembles or confidence regions around a learned dynamics model, which capture epistemic uncertainty induced by limited data support.
@@ 3 Methods
3.1 Robust regularized policy optimization
@@
Consequently, the induced optimal policy in Eq. (12) takes a Boltzmann form that is absolutely continuous with respect to the reference policy $\mu$. Moreover, the contraction factor remains $\gamma$ because the $\log$-$\exp$ mapping is non-expansive under $\|\cdot\|_{\infty}$ when $\alpha>0$, and the minimization over $p\in P$ preserves monotonicity. The regularization coefficient $\alpha$ controls the strength of regularization. Smaller $\alpha$ makes the update more greedy with respect to $Q^{*}$, while larger $\alpha$ yields a more conservative update that stays closer to $\mu$. This perspective helps explain why RRPI can be stabilized by keeping successive policies close in KL, while still allowing the algorithm to gradually move toward the soft-greedy solution implied by the robust regularized Bellman operator.
@@
Starting from any stochastic initial policy $\pi_{0}:S\to\Delta(A)$ with full support (i.e., $\pi_{0}(a|s)>0$ for all $(s,a)\in S\times A$), we generate a sequence $\{\pi_{i}\}_{i\geq 0}$ by repeatedly solving the robust sregularized ubproblem with the reference policy set to the previous iterate:
@@
Theorem 0.A.2 provides a formal link between optimizing the regularized surrogate $\widehat{\eta}(\pi,p,\mu)$ and improving the original objective $J(\pi)=\min_{p\in P}\eta(\pi,p)$. A key ingredient is the reference-policy update $\mu=\pi_{i}$, which makes each subproblem a local improvement step around the current iterate. This improvement step mirrors the role of trust regions while accounting for worst-case transition uncertainty through the inner minimization over $p\in P$. The requirement that $\pi_{0}$ has full support ensures that all subsequent iterates admit a well-defined regularization term $\log\frac{\pi}{\mu}$ (with $\mu=\pi_{i}$) and prevents irrevocably assigning zero probability to actions. This is important in robust settings, where an action that is suboptimal under nominal dynamics may become optimal under a different adversarial model in $P$. The ratio condition $\pi_{i}(a|s)/\pi_{i}(a^{\prime}|s)\to\infty$ whenever the limiting action-value of $a$ exceeds that of $a^{\prime}$ indicates that the sequence becomes increasingly greedy with respect to the limiting $Q$-function, thereby recovering an optimal robust policy of the original problem in the limit.
@@ 3.2 Implementations
Algorithm 1 summarizes the proposed procedure. We parameterize the action-value function $Q_{\theta}$ and the policy $\pi_{\phi}$ with neural networks. Constructing a suitable uncertainty set $P$ is nontrivial in practice: it must be rich enough to cover model errors while remaining computationally tractable for the inner minimization. In our implementation, we approximate $P$ using a learned dynamics model ensemble, where each transition model $p_{\nu_{i}}$ parameterizes a Gaussian distribution over next states and rewards [33, 8, 6]. Specifically, for each $(s,a)$, the ensemble induces a set of plausible next-state distributions and thus a tractable approximation of an uncertainty set. We instantiate the inner minimization by selecting, among the ensemble members, the transition model that yields the smallest one-step robust target in the Bellman backup. This worst-case member approximation serves as a computationally lightweight proxy to robust dynamic programming over $P$ and directly penalizes state-action pairs where the models disagree. Intuitively, when the dataset provides poor coverage, different ensemble members extrapolate differently, effectively enlarging the uncertainty set and causing the robust backup to propagate pessimistic value estimates. In contrast, in well-covered regions the ensemble predictions concentrate, and the robust backup reduces to near-nominal evaluation.
@@ 4 Experimental Results
In this section, we evaluate RRPI on standard offline RL benchmarks [2] with three complementary goals. First, we assess overall performance against strong baselines to verify the effectiveness of the proposed robust objective. Second, we analyze uncertainty estimation to examine whether the learned policy behaves conservatively in high-uncertainty regions, which is critical under distribution shift in offline settings. Third, we conduct ablation studies to isolate the contribution of the worst-case optimization and validate that robustness is essential for the gains.
@@ 4.1 Performance
We compare our method with a broad set of state-of-the-art baselines, including model-free algorithms such as CQL [10], DMG [17], and EPQ [14], as well as model-based approaches including MOReL [8], RAMBO [23], PMDB [6], and ADM [13]. As shown in Table 1, the proposed method achieves the best average performance on D4RL benchmarks [2]. Reference scores are taken from the original papers and [6] when available; for methods without reported numbers (e.g., ADM and DMG), we reproduce results by following the hyperparameter settings from the respective publications. In particular, compared with PMDB, our method outperforms it on 11 out of 18 environments and remains competitive on the remaining 7. This result suggests that robust optimization can provide better resilience under extreme transition perturbations than approaches that target percentile-based performance.
@@ 4.2 Uncertainty Estimation
Empirically, two patterns emerge: uncertainty spikes coincide with sharp drops in $Q(s,a)$, and the learned policy progressively avoids high-uncertainty regions, yielding a smoother $Q$-landscape. We attribute this phenomenon to the interaction between worst-case dynamics selection and regularized policy improvement. When the ensemble exhibits high disagreement at $(s,a)$, the uncertainty set contains plausible models that predict substantially different next states and rewards. The inner minimization then selects an adverse (yet plausible) transition within $P$, reducing the robust Bellman backup at $(s,a)$ and propagating pessimism to predecessor states. The subsequent KL-regularized policy update limits abrupt distributional shifts while still preferring actions with higher robust $Q$-values, causing the learned policy to progressively down-weight uncertain actions over iterations. In contrast to heuristic uncertainty penalties, the reduced values in high-disagreement regions here arise directly from optimizing a principled worst-case objective, providing a coherent explanation for the observed coupling between uncertainty spikes and $Q$-value drops.
@@ 5 Conclusions
Moving forward, we aim to tighten the theoretical-empirical gap in uncertainty estimation and explore the integration of multimodal observations (e.g., vision [25, 35, 38, 36]) into the RRPI framework to handle increasingly complex decision-making tasks.
@@ Appendix 0.A Proofs
Theorem 0.A.1
The robust regularized Bellman operator $\mathcal{T}$ is a $\gamma$-contraction mapping under the $\|\cdot\|_{\infty}$ norm. Starting from any initial action-value function $Q:S\times A\to\mathbb{R}$ and repeatedly applying $\mathcal{T}$, the resulting sequence converges to a unique fixed point $Q^{*}$. Moreover, the optimal robust policy for the surrogate objective $\widehat{\eta}(\pi,p,\mu)$ is given by
Proof
@@
Theorem 0.A.2
Starting from any stochastic initial policy $\pi_{0}:S\to\Delta(A)$ with full support (i.e., $\pi_{0}(a|s)>0$ for all $(s,a)\in S\times A$), we generate a sequence $\{\pi_{i}\}_{i\geq 0}$ by repeatedly solving the robust sregularized ubproblem with the reference policy set to the previous iterate:
Proof

### 14 Mar 2026 v2

@@ Abstract
Offline reinforcement learning (RL) enables data-efficient and safe policy learning without online exploration, but its performance often degrades under distribution shift. The learned policy may visit out-of-distribution state-action pairs where value estimates and learned dynamics are unreliable. To address policy-induced extrapolation and transition uncertainty in a unified framework, we formulate offline RL as robust policy optimization, treating the transition kernel as a decision variable within an uncertainty set and optimizing the policy against the worst-case dynamics. We propose Robust Regularized Policy Iteration (RRPI), which replaces the intractable max-min bilevel objective with a tractable KL-regularized surrogate and derives an efficient policy iteration procedure based on a robust regularized Bellman operator. We provide theoretical guarantees by showing that the proposed operator is a $\gamma$-contraction and that iteratively updating the surrogate yields monotonic improvement of the original robust objective with convergence. Experiments on D4RL benchmarks demonstrate that RRPI achieves strong average performance, outperforming recent baselines including percentile-based methods on the majority of environments while remaining competitive on the rest. Moreover, RRPI exhibits robust performance by aligning lower $Q$-values with high epistemic uncertainty, which prevents the policy from executing unreliable out-of-distribution actions.
@@ 1 Introduction
Reinforcement learning (RL) has become a key ingredient of modern decision-making systems, driving breakthroughs in complex domains such as strategic game playing, embodied intelligence, and large language model reasoning by optimizing long-horizon policies from interaction data [24, 14, 31]. Despite these successes, standard online RL typically relies on trial-and-error exploration, a process that can be prohibitively costly in high-stakes real-world applications [4, 8]. This has driven growing interest in offline RL [10], which aims to learn high-performing policies solely from pre-collected datasets without additional environment interaction.
@@
In this work, we pursue a robust policy optimization perspective. Instead of treating the learned dynamics as a point estimate, we regard the transition model as a decision variable that lies in an uncertainty set of plausible kernels, and we seek a policy that maximizes performance under the worst-case dynamics in this set [1]. This formulation incorporates dynamics uncertainty directly into the objective and yields a principled alternative to heuristic uncertainty penalties. Nevertheless, the induced max-min objective leads to a challenging bilevel optimization problem, whose exact solution is often computationally prohibitive in practical offline RL pipelines. To address this, we propose Robust Regularized Policy Iteration (RRPI), which introduces a KL-regularized surrogate objective and an associated robust regularized Bellman operator, enabling an efficient iterative algorithm with theoretical guarantees.
@@ 2.1 Related work
Offline RL learns a policy solely from a fixed dataset, and its core challenge arises from distributional shift: the learned policy may query state-action pairs that are poorly covered by the data, making value estimates unreliable [10, 19] and leading to extrapolation error [17]. Uncertainty in offline RL is often characterized as epistemic uncertainty induced by limited data support, which differs from aleatoric uncertainty originating from stochastic environment dynamics and observation noise [5, 3]. To mitigate this epistemic uncertainty, existing approaches primarily adopt conservative value learning that penalizes OOD actions [9, 29, 6, 28], regularized policy updates that constrain deviations from the behavior policy [26, 34], or explicit uncertainty estimation via ensembles or Bayesian methods [21, 5, 20]. In contrast to prior methods, we cast offline RL as a robust optimization problem over an uncertainty set, which incorporates dynamics uncertainty directly without requiring explicit uncertainty estimation. In this formulation, the dynamics model is treated as a variable rather than a fixed quantity. This method avoids heuristic penalties on OOD action values and does not require constraining the learned policy to remain close to the behavior policy. Building on this, we develop an iterative policy optimization algorithm with theoretical guarantees of convergence to an optimal solution.
@@ 2.2 Background
The modeling power and conservatism of robust MDPs depend critically on how $P$ is constructed. A common assumption is rectangularity, where $P$ factorizes across states or state-action pairs, enabling the inner minimization to decompose locally and making robust dynamic programming tractable. While convenient, rectangular sets can be overly conservative in high-dimensional problems since they allow independent worst-case perturbations at each $(s,a)$. In practice, $P$ is often built from function approximation, for example by using model ensembles, which capture epistemic uncertainty induced by limited coverage.
@@ 3 Methods
3.1 Robust regularized policy iteration
@@
Consequently, the induced optimal policy in Eq. (7) takes a Boltzmann form that is absolutely continuous with respect to the reference policy $\mu$. Moreover, the contraction factor remains $\gamma$ because the $\log$-$\exp$ mapping is non-expansive under $\|\cdot\|_{\infty}$ when $\alpha>0$, and the minimization over $p\in P$ preserves monotonicity. The coefficient $\alpha$ controls the strength of regularization. Smaller $\alpha$ makes the update more greedy with respect to $Q^{*}$, while larger $\alpha$ yields a more conservative update that stays closer to $\mu$. This perspective helps explain why RRPI can be stabilized by keeping successive policies close in KL, while still allowing the algorithm to gradually move toward the soft-greedy solution implied by the robust regularized Bellman operator.
@@
Starting from any stochastic initial policy $\pi_{0}:S\to\Delta(A)$ with full support (i.e., $\pi_{0}(a|s)>0$ for all $(s,a)\in S\times A$), we generate a sequence $\{\pi_{i}\}_{i\geq 0}$ by repeatedly solving the robust regularized subproblem with the reference policy set to the previous iterate:
@@
Theorem 3.2 provides a formal link between optimizing the regularized surrogate $\widehat{\eta}(\pi,p,\mu)$ and improving the original objective $J(\pi)=\min_{p\in P}\eta(\pi,p)$. A key ingredient is the reference-policy update $\mu=\pi_{i}$, which makes each subproblem a local improvement step around the current iterate. This improvement step mirrors the role of trust regions while accounting for worst-case transition uncertainty through the inner minimization over $p\in P$. The requirement that $\pi_{0}$ has full support ensures that all subsequent iterates admit a well-defined regularization term $\log\frac{\pi}{\mu}$ (with $\mu=\pi_{i}$) and prevents irrevocably assigning zero probability to actions. This is important in robust settings, where an action that is suboptimal under nominal dynamics may become optimal under a different adversarial model in $P$. The ratio condition $\pi_{i}(a|s)/\pi_{i}(a^{\prime}|s)\to\infty$ whenever the limiting action-value of $a$ exceeds that of $a^{\prime}$ indicates that the sequence becomes increasingly greedy with respect to the limiting $Q$-function, thereby recovering an optimal robust policy of the original problem in the limit.
@@ 3.2 Implementations
Algorithm 1 summarizes the proposed procedure. We parameterize the action-value function $Q_{\theta}$ and the policy $\pi_{\phi}$ with neural networks. Constructing a suitable uncertainty set $P$ is nontrivial in practice: it must be rich enough to cover model errors while remaining computationally tractable for the inner minimization. In our implementation, we approximate $P$ using a learned dynamics model ensemble, where each transition model $p_{\nu_{i}}$ parameterizes a Gaussian distribution over next states and rewards [30, 7, 5]. Specifically, for each $(s,a)$, the ensemble induces a set of plausible next-state distributions, yielding a tractable approximation of the uncertainty set. We instantiate the inner minimization by selecting, among the ensemble members, the transition model that yields the smallest one-step robust target in the Bellman backup. This worst-case member approximation serves as a computationally lightweight proxy to robust dynamic programming over $P$ and directly penalizes state-action pairs where the models disagree. Intuitively, when the dataset provides poor coverage, different ensemble members extrapolate differently, effectively enlarging the uncertainty set and causing the robust backup to propagate pessimistic value estimates. In contrast, in well-covered regions the ensemble predictions concentrate, and the robust backup reduces to near-nominal evaluation.
@@ 4 Experimental Results
In this section, we evaluate RRPI on D4RL benchmarks [2] with three goals. First, we assess overall performance against strong baselines to verify the effectiveness of the proposed robust objective. Second, we analyze uncertainty estimation to examine whether the learned policy behaves conservatively in high-uncertainty regions, which is critical under distribution shift in offline settings. Third, we conduct ablation studies to isolate the contribution of the worst-case optimization and validate that robustness is essential for the gains.
@@ 4.1 Performance
We compare our method with a broad set of state-of-the-art baselines, including model-free algorithms such as CQL [9], DMG [16], and EPQ [13], as well as model-based approaches including MOReL [7], RAMBO [21], PMDB [5], and ADM [12]. As shown in Table 1, our method achieves the best average performance on D4RL benchmarks. Reference scores are taken from the original papers; for methods without reported numbers (e.g., ADM and DMG), we reproduce results by following the hyperparameter settings from the respective publications. In particular, compared with PMDB, our method outperforms it on 11 out of 18 environments and remains competitive on the remaining 7. This result suggests that robust optimization can provide better resilience under extreme transition perturbations than approaches that target percentile-based performance.
@@ 4.2 Uncertainty Estimation
We observe that uncertainty spikes coincide with sharp drops in $Q(s,a)$, and the learned policy progressively avoids high-uncertainty regions, yielding a smoother $Q$-landscape. We attribute this phenomenon to the interaction between worst-case dynamics selection and regularized policy improvement. When the ensemble exhibits high disagreement at $(s,a)$, the uncertainty set contains plausible models that predict substantially different next states and rewards. The inner minimization then selects an adverse transition within $P$, reducing the robust Bellman backup at $(s,a)$ and propagating pessimism to predecessor states. The subsequent KL-regularized policy update limits abrupt distributional shifts while still preferring actions with higher robust $Q$-values, causing the learned policy to down-weight uncertain actions over iterations. Unlike heuristic uncertainty penalties, the reduced values in high-disagreement regions here arise directly from optimizing a principled worst-case objective, providing a coherent explanation for the observed coupling between uncertainty spikes and $Q$-value drops.
@@ 5 Conclusions
Moving forward, we aim to tighten the theoretical-empirical gap in uncertainty estimation and explore the integration of multimodal observations (e.g., vision [23, 32, 33]) into the RRPI framework to handle increasingly complex decision-making tasks.
@@ Appendix 0.A Proofs
0.A.1 Proof of Theorem 3.1
Proof(Proof of Theorem 3.1)
@@
0.A.2 Proof of Theorem 3.2
Proof(Proof of Theorem 3.2)